OpenStudy (anonymous):

if x=b+4 and y= b-3, then in terms of x and y b=?

OpenStudy (anonymous):

b = 4 - x b = 3 + y 2b=7 -x +y\[b=\frac{1}{2} (-x+y+7) \]